If this drug has a half-life of 8-hr instead of 4-hr, what is the amount of drug in the body at 24-hr and 48-hr?

time 0 hr 8 hr 16 hr 24 hr 32 hr 40 hr 48 hr
drug administered 100 mg 100 mg 100 mg 100 mg 100 mg 100 mg 100 mg
drug remaining 100 mg ? ? ? ? ? ?
Answer #1

Since Half life is 8 hr concentration will be 50 mg at 8hr and since you are administering 100 mg again it becomes 150 mg at 8 h

again at 16 h it will be 75 mg but since you are adding 100 mg it will be 175 mg

After another 8 h that is at 24 h it will be 175 mg/2 = 87.5 mg + 100 mg = 187.5 mg

at 32 h it will be 187.5/2 + 100 = 193.7 mg

At 40 h it will be 193.7/2 + 100 = 196.9 mg

at 48 h it will be 196.9/2 + 100 = 198.4 mg